lesie is 4 years older than john. the sum of their ages is 32. How old is Leslie? How old is John?
Answer by praseenakos@yahoo.com(507)

QUESTION:

lesie is 4 years older than john. the sum of their ages is 32. How old is Leslie? How old is John?
ANSWER:

Assume age of John is = x years.


Then age of Lesie = x + 4


Sum of their age is 32


==> x + x + 4 = 32


==> 2x + 4 = 32


Subtract 4 from both sides (Here we need to find the value of x....for that bring all other terms other than x)


==> 2x + 4 - 4 = 32 - 4


==> 2x = 28


Divide both sides of the equation by 2


==> 2x/2 = 28/2


==> x = 14


That age of John, x = 14 years.


Age of Lesie = x + 4 = 14 + 4 = 18 years.


To check your answer, it is given that sum of their age is 32....
here also 14 + 18 = 32

Hence the solution.



Hope you found this method useful


Regards.

Praseena.